TWO-DIMENSIONAL (2+n) REMPI SPECTROSCOPY: STATE INTERACTIONS, PHOTOFRAGMENTA-TIONS AND ENERGETICS OF THE HYDROGEN HALIDES

JINGMING LONG, HUASHENG WANG, \underline{AGUST KVARAN}, Science Institute, University of Iceland, Dunhagi 3, 107 Reykjavk, Iceland.

Mass spectra are recorded for one-colour (2 + n) resonance enhanced multiphoton ionization (REMPI) of HX (X = Cl ^{a}, Br) as a function of resonance excitation energy to obtain two-dimensional REMPI data. Perturbations due to Rydberg to ion-pair state interactions show as line shifts, ion signal intensity variations as well as band width broadenings depending on rotational quantum numbers, J ′. The data allow determination of parameters relevant to the nature and strength of state interactions as well as dissociation and ionization processes. Alterations in X^{+} and HX^{+} signal intensities prove to be very useful for spectra assignments.

aAgust Kvaran, Kristjan Matthiasson and Huasheng Wang, ” Two-Dimensional (2+n) Resonance Enhanced Multiphoton Ionization of HCl: State Interactions and Photorupture Channels via Low-Energy Triplet Rydberg States ”, J. Chem. Phys., \underline{131}(4), 044324, (2009).